# Optimized HDBSCAN clustering for reconstructing the merger history of the Milky Way: applications and limitations

### v1

@@ 4.1.1 Purity and completeness of clusters
Generally, HDBSCAN identifies clusters with a high-precision, which means that clusters are mostly composed of stars that belong to a single dominant progenitor. The lowest precision scores for the V-measure model are associated with specific physical scenarios. For Au7, the cluster linked to prog. 3210 (precision $P=0.58$) corresponds to the oldest major merger. In Au25, the cluster for prog. 692 ($P=0.62$) is a small stream dynamically similar to debris from the a recent major merger, prog. 736. For Au27,low-precision scores are found for clusters associated with prog. 9073 ($P=0.53$), prog 248 ($P=0.65$) and prog. 7219 ($P=0.55$). The former two progenitors are fragmented into three and two clusters, respectively; the low-precision clusters are also those that overlap with other debris in the phase space. The cluster associated to prog. 7219 is significantly contaminated by debris from the other two main merger events (prog. 168 and prog. 173), which slightly preceded the dominant progenitor and have not been recovered by the clustering model. The DBCV-optimized model identifies the same low-precision progenitors, with the addition of clusters for prog. 4643 in Au25, which also dynamically overlaps with debris from prog. 736.
On average, clusters associated with stellar streams show higher precision ($P=0.91$ for V-measure and $0.90$ for DBCV) than those associated with phase-mixed structures ($P=0.84$ for V-measure and $0.88$ for DBCV). These dynamically cold filamentary structures are less likely to overlap with debris from other progenitors than the phase-mixed structures, the latter often being contaminated by earlier mergers that remain undetected. This effect may also explain the higher precision found for clusters from the late mergers (that is, lookback time of infall $\leq 5$ Gyr) compared to earlier ones, a trend which is particularly evident for Au27. In terms of progenitor masses, clusters from the most massive mergers tend to have lower purity scores, which is likely due to the wider range in stellar properties (features) of their debris.
@@
In contrast to the high-precision scores, the recall scores are significantly lower for both models. The average recall ($R$) values for clusters of stellar streams are 0.40 (for V-measure) and 0.24 (for DBCV), while for clusters of phase-mixed structures are 0.16 (for V-measure) and 0.13 (for DBCV), respectively. This is primarily due to two factors: a large number of particles are being classified as noise, and the fragmentation of single progenitors into multiple clusters. The latter effect explains the lower recall of the DBCV-optimized model, which tends to produce more clusters per progenitor. For example, for prog. 2197 in Au7, $85\%$ of its particles are classified as noise in both models, while $10\%$ are split across six separate clusters.
The fragmentation of the debris of a progenitor into multiple clusters is expected for the most massive progenitors. Those are typically accreted late and experience a prolonged chemical evolution; therefore, they tend to consist of multiple stellar populations and their debris spans a wide range in chrono-chemodynamical properties. To investigate this hypothesis, we calculated the total variance of each progenitor’s debris in the 12-dimensional parameter space (Fig. 5). Progenitors that are split into several clusters generally show higher variance than the rest. However, a large variance does not always lead to fragmentation. In cases where the debris are too widely dispersed, the algorithm either assigns many of those stars to other clusters or classifies them as noise. Stars are most often classified as noise when debris from different progenitors overlap in the parameter space. This effect is illustrated in Fig. 6, where progenitors with the largest number of stars flagged as noise are also those with the smallest Mahalanobis distance to the progenitor that dominates the noise population.
@@ 4.1.2 The assembly histories inferred from clustering
We further investigate the assembly histories of the three galaxies, as represented by their merger trees constructed from the clusters recovered by the optimized algorithm. The full merger trees for the sample galaxies, computed directly from the simulations with the methods described in Section 2, are shown in the top panel of Fig. 7. Each branch in the tree represents a progenitor galaxy that joins the main branch (shown in black) when the progenitor infalls onto the host. The thickness of each branch is proportional to the number of accreted particles that the progenitor satellite contributes to the host at $z=0$. The dashed lines indicate progenitors that contribute less than 1% of the total number of accreted particles at $z=0$. The middle and bottom panels show the merger trees as inferred from the clustering results obtained with the V-measure-optimized and DBCV-optimized HDBSCAN models, respectively. Here, a branch is included only if its corresponding progenitor is dominant in at least one identified cluster.
@@
As expected, both models are more successful in recovering more massive mergers (those contributing >1% of the total accreted particles, shown as solid lines and reported in Table 1). The major merger events that are missed are predominantly those that occur at high redshift ($z_{\rm acc}\gtrsim 2$), with the exception of prog. 6355 (light green) in Au27. This progenitor merged with the main galaxy around the same time as prog. 205, a satellite about ten times as massive; thus, its stellar debris was likely masked in the feature space by the dominant signature of the larger satellite, preventing its identification by the clustering algorithm.
@@ 4.2 Clustering in accreted + in situ stellar haloes
Fig. 9 shows the Mahalanobis distance between the 12-D distributions of the debris from the progenitors and the in situ population. This shows that the oldest major mergers (e.g., prog. 3210 in Au7, prog. 41397 in Au25 and prog. 7219 and in Au27) have the largest overlap with the in situ population in the clustering space. Therefore, the particles from these mergers tend to be associated with the same cluster of the in situ population. A stricter density requirement (i.e., a larger value of min_samples) causes a fragmentation of the sparser regions of the in situ cluster. Because a larger number of progenitors in Au7 are mixed with the in situ ones than in the other galaxies, the fragmentation of the in situ clusters leads to the formation of groups dominated by accreted particles, hence increasing the overall V-measure score of the clustering configuration in spite of a decrease in the cluster density.
@@ 4.2.1 Purity and completeness of clusters
As shown in Fig. 10, the contamination from the in situ particles only slightly degrades the precision of the identified accreted clusters, with most still being composed of well over 50% of stars from their dominant progenitor. The average precision for clusters associated with stellar streams and phase-mixed substructures is 0.91 and 0.81, respectively, for the V-measure optimized model, and 0.87 and 0.83 for the DBCV one. Conversely, the recall scores decreased on average: 0.23 (V-measure) and 0.27 (DBCV) for streams, 0.03 (V-measure) and 0.04 (DBCV) for phase-mixed structures. This is due to increased fragmentation, as shown in Fig. 11. A notable exception is Au25, whose larger progenitors, which appear as stellar streams, are recovered with similar or even higher recall scores than in the accreted-only analysis. This counterintuitive result likely arises from a contrast effect: the introduction of the dense, clustered in situ population makes the relatively less dense but coherent stellar streams stand out more clearly, allowing the algorithm to capture a larger fraction of their members. The in situ dominated clusters are also particularly pure and have high recall scores. This is probably due to both the efficiency of the algorithm of isolating the dense in situ component into one or a few coherent groups, and the significantly larger number of in situ particles compared to the accreted ones in these regions of the clustering space. The “inferred” lookback infall time for the in situ clusters, i.e., the $90^{\rm th}$ percentiles of the cluster age distributions, appears to vary significantly, which indicates that particles that formed in different star formation episodes also have different chrono-chemodynamical properties. In all galaxies, the only clusters with an associated lookback infall time >10>10 Gyr are in situ dominated.
@@ 4.2.2 The assembly histories inferred from clustering
A comparison of the results with and without in situ contamination reveals that the primary impact of the in situ particles is not a degradation in the purity of recovered clusters, but rather the complete non-detection of certain progenitor groups. To better understand which mergers are more likely to be affected, we analyze the merger tree plots in Fig. 12. Similarly to Fig. 7, the actual merger history of the galaxies is shown in the top panel, while the middle and bottom panels only display those merger events which have been identified as the dominant progenitors of the HDBSCAN clusters. Again, the continuous lines indicate merger events that contribute at least 1% of the total accreted stars in the galaxy at $z=0$, and the thickness of the line is proportional to the actual number.
@@
Figure 12: Merger trees for the galaxies in the sample as reconstructed by the V-measure (middle) and DBCV (bottom) optimized HDBSCAN models when contamination from in situ stars is included. Top panel shows the full merger trees. The correspondence between lines and progenitors is the same as described in Fig. 7 for the accreted-only case.
@@ 5 Discussion
Our analysis shows that an optimized HDBSCAN model can produce pure clusters of accreted stars even in environments strongly contaminated by the in situ population (see Fig. 10). This is in contrast with the results of Thomas et al. (2025) who use the single-linkage clustering methodology developed by Lövdal et al. (2022) and find that it predominantly returns clusters composed in majority by in situ stars or by a combination of debris from different progenitors (see Section 4.1 in Thomas et al. 2025). We believe that the optimization of the clustering algorithm is the crucial factor in correctly identifying the clusters made of accreted stars. Also, while the Lövdal et al. (2022) clustering algorithm uses only properties related to IoM, our study uses an extended, 12-D parameter (feature) space that combines chemistry, kinematics and ages. There are, of course, other factors that may explain the discrepancies between the results of the two methods. The Lövdal et al. 2022 methodology imposes the creation of “background haloes” in order to determine the significance score of the clusters, which may introduce artificial density variations in the clustering space. There are also differences between the two Auriga data sets in terms of their numerical resolution, as Thomas et al. (2025) use the “level 3” suite, which are higher in resolution than the “level 4” suite used in our analysis. Moreover, although both studies use samples of star particles with similar halo-type kinematics, Thomas et al. (2025) focus on solar-neighbourhood regions, whereas we study the entire haloes.
@@
Our optimized HDBSCAN model can retrieve mergers up to $z_{\rm acc}\sim 3$ (see Fig. 7). However, when applied to a data set contaminated by in situ stars, only galaxies accreted after $z_{\rm acc}\simeq 1$ are retrieved as dominant progenitors of the clusters. This is similar to the results of Thomas et al. (2025), which show that their methodology is effective for mergers that occur in the last 676-7 Gyr. However, because the purity of the HDBSCAN clusters does not change significantly between the accreted-only and accreted + in situ data sets, we argue that the presence of the in situ population in the clustering data set mainly impacts the number and type of detectable mergers. We note that the problem could be attenuated when using observations, as the Auriga simulated haloes may have a higher fraction of in situ stars than measured in the Milky Way (Monachesi et al., 2019). Therefore, we expect that a higher fraction of the assembly history of the Milky Way could be retrieved with our methodology than implied by the results from clustering on accreted + in situ stars in Section 4.2. Moreover, there are other machine learning methods that can very efficiently separate in situ stars from accreted stars (Ostdiek et al., 2020; Sante et al., 2024), before applying any clustering methods. Using these methods in combination, one would expect a higher rate of success in finding substructures, even in the case of heavy contamination by in situ stars.
@@
As already mentioned, in our analysis clusters are obtained from the full chrono-chemodynamical properties of the accreted particles. However, replicating this parameter space for stars in the Milky Way might be challenging. Although kinematics measurements are accurately provided for tens of millions of stars by Gaia and self-consistent chemical abundances are becoming available for a growing number of stars through the combination of data from different spectroscopic surveys (Thomas et al., 2024; Horta et al., 2025), stellar ages remain accessible only for a fraction of them. Moreover, estimates of the total energy of stars in the Milky Way rely on assumptions on the gravitational potential of the Galaxy (e.g., McMillan, 2017). The methodology presented in this paper does not account for any uncertainty or incompleteness in the stellar properties. Hence, more studies are needed to understand the implications of this missing information on the capability of HDBSCAN to recover past mergers.

### v2

@@ 4.1.1 Purity and completeness of clusters
Generally, HDBSCAN identifies clusters with a high-precision, which means that clusters are mostly composed of stars that belong to a single dominant progenitor. The lowest precision scores for the V-measure model are associated with specific physical scenarios. For Au7, the cluster linked to prog. 3210 (precision $P=0.58$) corresponds to the oldest major merger. In Au25, the cluster for prog. 692 ($P=0.62$) is a small stream dynamically similar to debris from a recent major merger, prog. 736. For Au27,low-precision scores are found for clusters associated with prog. 9073 ($P=0.53$), prog 248 ($P=0.65$) and prog. 7219 ($P=0.55$). The former two progenitors are fragmented into three and two clusters, respectively; the low-precision clusters are also those that overlap with other debris in the phase space. The cluster associated with prog. 7219 is significantly contaminated by debris from the other two main merger events (prog. 168 and prog. 173), which slightly preceded the dominant progenitor and have not been recovered by the clustering model. The DBCV-optimized model identifies the same low-precision progenitors, with the addition of clusters for prog. 4643 in Au25, which also dynamically overlaps with debris from prog. 736.
On average, clusters associated with stellar streams show higher precision ($P=0.91$ for V-measure and $0.90$ for DBCV) than those associated with phase-mixed structures ($P=0.84$ for V-measure and $0.88$ for DBCV). These dynamically cold filamentary structures are less likely to overlap with debris from other progenitors than the phase-mixed structures, the latter often being contaminated by earlier mergers that remain undetected. This effect may also explain the higher precision found for clusters from the late mergers (that is, lookback time of infall $\leq 5$ Gyr) compared to earlier ones; this effect is particularly evident for Au27 (see the rightmost subpanel in top panel of Fig. 3). In terms of progenitor masses, clusters from the most massive mergers tend to have lower purity scores, which is likely due to the wider range in stellar properties (features) of their debris.
@@
In contrast to the high-precision scores, the recall scores are significantly lower for both models. The average recall ($R$) values for clusters of stellar streams are 0.40 (for V-measure) and 0.24 (for DBCV), while for clusters of phase-mixed structures are 0.16 (for V-measure) and 0.13 (for DBCV), respectively. This is primarily due to two factors: a large number of particles are being classified as noise, and the fragmentation of single progenitors into multiple clusters. The latter effect explains the lower recall of the DBCV-optimized model, which tends to produce more clusters per progenitor. For example, for prog. 2197 in Au7, $85\%$ of its particles are classified as noise in both models, while $10\%$ are split into six separate clusters.
The fragmentation of the debris of a progenitor into multiple clusters is expected for the most massive progenitors. Those are typically accreted late and experience a prolonged chemical evolution; therefore, they tend to consist of multiple stellar populations and their debris spans a wide range in chrono-chemodynamical properties. To investigate this hypothesis, we calculated the total variance of the debris of each progenitor in the 12-dimensional parameter space (Fig. 5). Progenitors that are split into several clusters generally show higher variance than the rest. However, a large variance does not always lead to fragmentation. In cases where the debris are too widely dispersed, the algorithm either assigns many of those stars to other clusters or classifies them as noise. Stars are most often classified as noise when debris from different progenitors overlap in the parameter space. This effect is illustrated in Fig. 6, where the progenitors with the largest number of stars flagged as noise are also those with the smallest Mahalanobis distance from the progenitor that dominates the noise population.
@@ 4.1.2 The assembly histories inferred from clustering
We further investigate the assembly histories of the three galaxies, as represented by their merger trees constructed from the clusters recovered by the optimized algorithm. The full merger trees for the sample galaxies, computed directly from the simulations with the methods described in Section 2, are shown in the top panel of Fig. 7. Each branch in the tree represents a progenitor galaxy that joins the main branch (shown in black) when the progenitor falls into the host. The thickness of each branch is proportional to the number of accreted particles that the progenitor satellite contributes to the host at $z=0$. The dashed lines indicate progenitors that contribute less than 1% of the total number of accreted particles at $z=0$. The middle and bottom panels show the merger trees as inferred from the clustering results obtained with the V-measure-optimized and DBCV-optimized HDBSCAN models, respectively. Here, a branch is included only if its corresponding progenitor is dominant in at least one identified cluster.
@@
As expected, both models are more successful in recovering more massive mergers (those contributing >1% of all accreted particles, shown as solid lines and reported in Table 1). The major merger events that are missed are predominantly those that occur at high redshift ($z_{\rm acc}\gtrsim 2$), with the exception of prog. 6355 (light green) in Au27. This progenitor merged with the main galaxy around the same time as prog. 205, a satellite about ten times as massive; thus, its stellar debris was likely masked in the feature space by the dominant signature of the larger satellite, preventing its identification by the clustering algorithm.
@@ 4.2 Clustering in accreted + in situ stellar haloes
Fig. 9 shows the Mahalanobis distance between the 12-D distributions of the debris from the progenitors and the in situ population. This shows that the oldest major mergers (e.g., prog. 3210 in Au7, prog. 41397 in Au25 and prog. 7219 and in Au27) have the largest overlap with the in situ population in the clustering space. Therefore, the particles from these mergers tend to be associated with the same cluster of the in situ population. A stricter density requirement (i.e., a higher value of min_samples) causes a fragmentation of the sparser regions of the in situ cluster. Because a larger number of progenitors in Au7 are mixed with the in situ ones than in the other galaxies, the fragmentation of the in situ clusters leads to the formation of groups dominated by accreted particles, hence increasing the overall V-measure score of the clustering configuration despite of a decrease in the cluster density.
@@ 4.2.1 Purity and completeness of clusters
As shown in Fig. 10, the contamination from in situ particles only slightly degrades the precision of the identified accreted clusters, with most still being composed of well over 50% of stars from their dominant progenitor. The average precision for clusters associated with stellar streams and phase-mixed substructures is 0.91 and 0.81, respectively, for the V-measure optimized model, and 0.87 and 0.83 for the DBCV one. Conversely, the recall scores decreased on average: 0.23 (V-measure) and 0.27 (DBCV) for streams, 0.03 (V-measure) and 0.04 (DBCV) for phase-mixed structures. This is due to increased fragmentation, as shown in Fig. 11. A notable exception is Au25, whose larger progenitors, which appear as stellar streams, are recovered with similar or even higher recall scores than in the accreted-only analysis. This counterintuitive result likely arises from a contrast effect: the introduction of the dense, clustered in situ population makes the relatively less dense but coherent stellar streams stand out more clearly, allowing the algorithm to capture a larger fraction of their members. The in situ dominated clusters are also particularly pure and have high recall scores. This is probably due to both the efficiency of the algorithm of isolating the dense in situ component into one or a few coherent groups, and the significantly larger number of in situ particles compared to the accreted ones in these regions of the clustering space. The “inferred” lookback infall time for the in situ clusters, i.e., the $90^{\rm th}$ percentiles of the cluster age distributions, appears to vary significantly, which indicates that particles that formed in different star formation episodes also have different chrono-chemodynamical properties. In all galaxies, the only clusters with an associated lookback infall time >10>10 Gyr are in situ dominated.
@@ 4.2.2 The assembly histories inferred from clustering
A comparison of the results with and without in situ contamination reveals that the primary impact of the in situ particles is not a degradation in the purity of recovered clusters, but rather the complete non-detection of certain progenitor groups. To better understand which mergers are more likely to be affected, we analyze the merger tree plots in Fig. 12. Similarly to Fig. 7, the actual merger history of the galaxies is shown in the top panel, while the middle and bottom panels only display those merger events which have been identified as the dominant progenitors of the HDBSCAN clusters. Again, the continuous lines indicate merger events that contribute at least 1% of the total accreted stars in the galaxy (after the velocity selection cut) at $z=0$, and the thickness of the line is proportional to the actual number.
@@
Figure 12: Merger trees for the galaxies in the sample as reconstructed by the V-measure (middle) and DBCV (bottom) optimized HDBSCAN models when contamination from in situ stars is included. Top panel shows the full merger trees. The correspondence between lines and progenitors is the same as described in Fig. 7 for the accreted-only case. The style and thickness of the lines reflects the new fraction of accreted stars from each progenitor after the velocity selection cut is applied.
4.3 The impact of stellar ages
As already mentioned, in our analysis clusters are obtained from the full chrono-chemodynamical properties of the accreted particles. However, replicating this parameter space for stars in the Milky Way might be challenging. Although kinematics measurements are accurately provided for tens of millions of stars by Gaia and self-consistent chemical abundances are becoming available for a growing number of stars through the combination of data from different spectroscopic surveys (Thomas et al., 2024; Horta et al., 2025), stellar ages remain accessible only for a fraction of them. Therefore, a more realistic application of our methodology would involve only the dynamical and chemical properties of the stars in the Milky Way. In order to assess the quality of the merger tree reconstruction under these conditions, we repeat the analysis discussed in Sections 4.1 and 4.2 excluding the ages of the star particles ($\tau_{\text{form}}$) from the vector of stellar properties describing the clustering space. For the sake of conciseness, we consider only Au27, which is the galaxy with the assembly history most similar to the Milky Way in our sample, for this part of the analysis.
Similarly to Figs. 2 and 8, the distribution in the $E-L_{z}$ plane of the star particles in Au27 coloured based on their HDBSCAN group are reported in Fig. 13 for both the clustering considering only accreted stars (left plots) and including in-situ contamination (right plots). In both cases, the HDBSCAN parameters have been fine-tuned to adapt to the new clustering space without the age information. The top plots show the results of the clustering model optimised using the V-measure metric, while the bottom ones refer to the optimisation obtained with the DBCV metric. By comparing Fig. 13 with Figs. 2 and 8, it can be noticed that excluding the stellar ages results in a larger number of clusters. The clustering model optimised with V-measure returns 30 and 38 clusters in this scenario for the accreted and accreted + in-situ datasets, while only 11 and 15 clusters, respectively, are identified by the same model when the age information is provided. Similarly, the DBCV-optimised model returns a larger number of clusters for both the accreted-only (17 vs 16) and accreted + in-situ (46 vs 20) cases. This effect can be explained as fragmentation of the progenitors identified in the analysis with stellar ages into a larger number of clusters and a consequent decrease in the average completeness of the retrieved groups in both the accreted-only (0.15 vs 0.29, V-measure and 0.24 vs 0.19, DBCV) and accreted + in-situ (0.08 vs 0.16, V-measure and 0.09 vs 0.15, DBCV) datasets. The average purity of the groups is similar whether the clustering is performed with (0.78, V-measure, and 0.86, DBCV, for the accreted-only stars and 0.88, V-measure, and 0.87, DBCV, for the accreted + in-situ stars) or without the age (0.90, V-measure, and 0.84, DBCV, for the accreted-only stars and 0.87, V-measure, and 0.87, DBCV, for the accreted + in-situ stars) of the star particles. Therefore, the identification of pure samples of stars accreted from a single progenitors is still possible without the knowledge of their formation time. However, the inclusion of such information is important for grouping the accreted stars in a number of substructures more representative to the number of actual progenitor galaxies. Therefore, if stellar ages cannot be determined, the development of a methodology for further grouping the star samples identified by HDBSCAN is crucial for a correct inference of the Galaxy assembly history.
Figure 13: Density distribution in the $E-L_{z}$ plane for the stars in Au27 colour-coded by their HDBSCAN label. The clustering analysis is repeated excluding the age of the star particles from both the accreted-only and accreted + in-situ datasets. The plot format follows Figs. 2 and 8.
@@ 5 Discussion
Our analysis shows that an optimized HDBSCAN model can produce pure clusters of accreted stars even in environments strongly contaminated by the in situ population (see Fig. 10). This is in contrast with the results of Thomas et al. (2025) who use the single-linkage clustering methodology developed by Lövdal et al. (2022) and find that it predominantly returns clusters composed in majority by in situ stars or by a combination of debris from different progenitors (see Section 4.1 in Thomas et al. 2025). We believe that the optimization of the clustering algorithm is the crucial factor in correctly identifying the clusters made of accreted stars. Also, while the Lövdal et al. (2022) clustering algorithm uses only properties related to IoM, our study uses an extended, 12-D parameter (feature) space that combines chemistry, kinematics and ages. There are, of course, other factors that may explain the discrepancies between the results of the two methods. The Lövdal et al. 2022 methodology imposes the creation of “background haloes” in order to determine the significance score of the clusters, which may introduce artificial density variations in the clustering space. There are also differences between the two Auriga data sets in terms of their numerical resolution, as Thomas et al. (2025) use the “level 3” suite, which are higher in resolution than the “level 4” suite used in our analysis. As the stellar masses of galaxies, both in situ and accreted, increase at higher resolution (see e.g., Grand et al. 2021 and Riley et al. 2025), it is possible that the contamination by in situ stars in the sample also increases. Moreover, although both studies use samples of star particles with similar halo-type kinematics, Thomas et al. (2025) focus on solar neighbourhood regions, whereas we study the entire haloes.
Our optimized HDBSCAN model can retrieve mergers up to $z_{\rm acc}\sim 3$ (see Fig. 7). However, when applied to a data set contaminated by in situ stars, only galaxies accreted after $z_{\rm acc}\simeq 1$ are retrieved as dominant progenitors of the clusters. From an observational point of view, this suggests that accretion events >10 Gyr old may not be found by clustering methods, or that any clusters from this type of events are likely dominated by in situ material. This suggests that observational samples representing early accretions in the Galaxy, such as GES (Helmi et al., 2018; Belokurov et al., 2018), Kraken (Kruijssen et al., 2019), or Shiva & Shakti (Malhan and Rix, 2024) could be significantly contaminated by in situ stars.
@@
These findings are similar to the results of Thomas et al. (2025), who show that their methodology is effective for mergers that occurred in the last 676-7 Gyr. However, because the purity of the HDBSCAN clusters does not change significantly between the accreted-only and accreted + in situ data sets, we argue that the presence of the in situ population in the clustering data set mainly impacts the number and type of detectable mergers. We note that the problem could be attenuated when using observations, as the Auriga simulated haloes may have a higher fraction of in situ stars than measured in the Milky Way (Monachesi et al., 2019). Therefore, we expect that a higher fraction of the assembly history of the Milky Way could be retrieved with our methodology than implied by the results from clustering on accreted + in situ stars in Section 4.2. Moreover, there are other machine learning methods that can very efficiently separate in situ stars from accreted stars (Ostdiek et al., 2020; Sante et al., 2024), before applying any clustering methods. Using these methods in combination, one would expect a higher rate of success in finding substructures, even in the case of heavy contamination by in situ stars.
@@
In Sections 4.14.2, the clusters were obtained from the full (12-D) chrono-chemodynamical properties of the accreted particles. However, replicating this 12-D parameter space for stars in the Milky Way may be challenging. Although the kinematical measurements are accurately provided for tens of millions of stars by Gaia and self-consistent chemical abundances are becoming available for a growing number of stars through the combination of data from different spectroscopic surveys (Thomas et al., 2024; Horta et al., 2025), stellar ages remain accessible only for a fraction of them. In Section 4.3 we examined a more realistic scenario where the ages were not available in the set of features. We found that, on average, the purity of the clusters was not affected, but it was a slight decrease in their completeness. Overall, the cluster became ever more fragmented.
Furthermore, the methodology presented in this paper does not account for any uncertainty or incompleteness in the stellar properties. Hence, more studies are needed to understand the implications of this missing information on the capability of HDBSCAN to recover past mergers.